OpenStudy (jocelynstalnaker):

I need help in for number four. Willing to give metal(:

OpenStudy (jocelynstalnaker):

A company produces accessories for smart phones and tablets. The profit on each smart phone case is $2, and the profit on each tablet case is $3. The company made a profit of $1,200 on the cases last month. The equation 2x + 3y = 1,200 represents the company's profit from cases last month, where x is the number of smart phone cases sold and y is the number of tablet cases sold.

OpenStudy (jocelynstalnaker):

3.Write the equation in function notation. Explain what the graph of the function represents. Be sure to use complete sentences. 4.Graph the function using one of the following two options below. On the graph, make sure to label the intercepts. -You may graph your equation by hand on a piece of paper and scan your work. -You may graph your equation using graphic technology that can be found in the Course Information area.

OpenStudy (jocelynstalnaker):

I did number three, which is F(x) = 2x + 1200

OpenStudy (jocelynstalnaker):

I just need to know how to get the points.

OpenStudy (jdoe0001):

F(x) = 2x + 1200 <--- how did you get it?

OpenStudy (yanasidlinskiy):

Is this what you were looking for??

OpenStudy (jocelynstalnaker):

Yes, @YanaSidlinskiy

OpenStudy (jocelynstalnaker):

A company produces accessories for smart phones and tablets. The profit on each smart phone case is $2, and the profit on each tablet case is $3. The company made a profit of $1,200 on the cases last month. The equation 2x + 3y = 1,200 represents the company's profit from cases last month, where x is the number of smart phone cases sold and y is the number of tablet cases sold.

OpenStudy (jocelynstalnaker):

@jdoe0001

OpenStudy (yanasidlinskiy):

Is there a 1 and 2 for this??

OpenStudy (jocelynstalnaker):

1.Change the equation into slope-intercept form. Identify the slope and y-intercept of the equation. Be sure to show all of your work. 2.Describe how you would graph this line using the slope-intercept method. Be sure to write in complete sentences. 3.Write the equation in function notation. Explain what the graph of the function represents. Be sure to use complete sentences. 4.Graph the function using one of the following two options below. On the graph, make sure to label the intercepts. -You may graph your equation by hand on a piece of paper and scan your work. -You may graph your equation using graphic technology that can be found in the Course Information area. 5.Suppose in the next month, the total profit on smart phone cases and tablet cases is $1,500. The profit amounts are the same, $2 for smart phone case and $3 for the tablet case. In a paragraph of at least three sentences, explain how the graphs of the functions for the two months are the same and how they are different. Be sure to use complete sentences. 6.Below is a graph that represents the total profits for a third month. Write the equation of the line that represents this graph. Show your work or explain how you determined the equations.

OpenStudy (jdoe0001):

ahemm I read that revise number 3

OpenStudy (jocelynstalnaker):

I have already done 1-3 and 5 and 6. Just not 4

OpenStudy (jocelynstalnaker):

Its not right? @jdoe0001

OpenStudy (jdoe0001):

nope as I read it

OpenStudy (jdoe0001):

3.Write the equation in function notation <---- they don't quite say rewrite is using f(x) only

OpenStudy (jocelynstalnaker):

Oh.. Then can you help me? I dont quite understand.

OpenStudy (jocelynstalnaker):

Understand it*

OpenStudy (jdoe0001):

f(x) = y thus 2x + 3y = 1,200 solve for "y" then replace y for f(x)

OpenStudy (jocelynstalnaker):

2x + 3y = 1200 2x-1200=-3y 1/3*(2x-1200)=1/3*-3y 2/3x-400=-y -2/3x+400=y

OpenStudy (jocelynstalnaker):

SO f(x)=-2/3x+400

OpenStudy (jocelynstalnaker):

Right?

OpenStudy (jdoe0001):

well.. yes

OpenStudy (jdoe0001):

-2/3x+400=y y= f(x) = -2/3x+400

OpenStudy (jocelynstalnaker):

Oh ok! But now how would i put it on a graph?

OpenStudy (jdoe0001):

or you can just pick a couple of random "x" values, get "y" and plot the line is a line and for that all you need is 2 points anyway

OpenStudy (jocelynstalnaker):

Thank you so much!

OpenStudy (jdoe0001):

yw